The position of a particle moving in a straight line is given by x = 3 + 4t + 3t2, where x is in
metre and time is in second. Find the values of the following physical quantities for the particle
at t = 2s.
1
Expert's answer
2017-12-12T12:49:06-0500
The particle motion law is given by
x = 3 + 4t + 3t^2
At the moment of time t=2 s
a) Position
x=3+4×2+3×2^2=23 m
b) Velocity
v(t)=x' (t)=4+6t
v(2)=4+6×2=16 m/s
c) Acceleration
a(t)=x'' (t)=6 m/s^2